Bibliography

thesis

23/03-04 16:19

[akselsen1991a] Sigmund Akselsen. Mediating Knowledge — Case Study: A Computer-Based Learning Environment for Education and Training of Fish-Health Personel. Dr. scient thesis, Department of Computer Science, University of Tromsø, March 1991.

[andreassen1992a] Jan Andreassen. A datalogger with telemetric datatransfer. Semester thesis, Department of Computer Science, University of Tromsø, June 1992.

[andreassen1992b] Jan Andreassen. A datalogger with packet radio and satellite communication interface. Diploma thesis, Department of Computer Science, University of Tromsø, December 1992.

[bates1994a] John Bates. Presentation Support for Distributed Multimedia Applications. PhD thesis, Computer Laboratory, University of Cambridge, January 1994.

[blair1994c] Lynne Blair. The Formal Specification and Verification of Distributed Multimedia Systems. Dr. thesis, Department of Computer Science, Lancaster University, September 1994.

[blum1997b] Christian Blum. Platform Support for the Development and Deployment of Multipoint Multimedia Application. PhD thesis, École Polytechnigue Fédérale de Lusanne, October 1997.

[bogdan1990a] Eniko Bogdan. Turtle graphics on a functional platform. Master’s thesis, the University of Wisconsin-Milwaukee, April 1990.

[brattli1996a] Dag Brattli. The software network: Providing continuos network connectivity for multihoming mobile computers. Master of engineering thesis, Department of Computer Science, Institute for Mathamatical and Physical Science, University of Tromsø, December 1996.

[cao1997a] Wen Cao. Achieving Efficient Cooperation in a Multi-Agent World: the Twin-Base Agent Modeling Approach. Doctoral dissertation, Department of Computer Science, Faculty of Science, University of Tromsø, August 1997.

[cheong1992a] Fah-Chun Cheong. OASIS: An Agent-Oriented Programming Language for Heterogeneous Distributed Environment. PhD thesis, The University of Michigan, 1992.

[costa2001a] Fabio Costa. Combining Meta-Information Management and Reflection in an Architecture for Configurable and Reconfigurable Middleware. PhD thesis, Computing Department, Lancaster University, August 2001.

[coulson1993c] Geoffrey Coulson. Multimedia Application Support in Opend Distributed Systems. PhD thesis, Computing Department, Lancaster University, April 1993.

[crosby1995a] Simon Andrew Crosby. Performance Managment in ATM Networks. PhD thesis, St. John’s Colleg, University of Cambridge, May 1995.

[danner1996a] G. Danner. Multimedia transport services. Diploma-thesis, Lehrstuhl für Praktische Informatik IV, Universität Mannheim / Computing Department, Lancaster University, 1996.

[diesen1995a] Dag Diesen. Specyfying Algorithms Using Evolving Algebra. Implementation of Functional Programming Languages. Dr. scient thesis, Department of Informatics, University of Oslo, March 1995.

[elde1999a] Morten Elde. Nuclear encoded homing endonucleases from Didymium and Naegleria. Dr. thesis, Department of Molecular Cell Biology, Institute of Medical Biology, University of Tromsø, September 1999.

[elstad1994a] Kristian Elstad. Bilde komprimering ved bruk av Wavelet transform koding. Diplomoppgave i informatikk, Seksjon for Informatikk, Institutt for Matematiske Realfag, Universitetet i Tromsø, June 1994.

[gancedo1999a] David Sánches Gancedo. QoSMonAuTA, QoS monitoring and adaptation using timed automata. Master’s thesis, Lancaster University, UK, September 1999.

[guttag1975a] John Guttag. The Specification and Application to Programming of Abstract Data Types. PhD thesis, University of Toronto, August 1975.

[hanssen1994a] Øyvind Hanssen. Objekt-adapter og programmeringsgrensesnitt for ANSAware applikasjoner. Hovedoppgave i informatikk, Seksjon for Informatikk, Institutt for Matematiske Realfag, Universitetet i Tromsø, 1994.

[hartvigsen1992a] Gunnar Hartvigsen. Domain limitations and consequences of domain limitations on the system role in knowledge-based systems within financial analysis and strategic market planning. Dr. scient thesis, Department of Computer Science, University of Tromsø, June 1992.

[hayden1998a] Mark Garland Hayden. The Ensemble System. PhD thesis, Faculty of the Graduate School of Cornell University, January 1998.

[hayter1993a] Mark David Hayter. A Workstation Architecture to Support Multimedia. PhD thesis, St. John’s College, University of Cambridge, September 1993.

[helander1994a] Johannes Helander. Unix under Mach — the Lites server. Master’s thesis, Helsinki University of Technology, Faculty of Information Technology, Department of Computer Science, December 1994.

[helme1989a] Arne Helme. Feiltoleranse i en distribuert applikasjon. Cand. mag. oppgave, Universitetet i Tromsø, 1989.

[helme1992a] Arne Helme. Tildeling av prosessorer til prosesser i et distribuert system ved anvendelse av en multidimensjonal algoritme. Cand. scient oppgave, Universitetet i Tromsø, August 1992.

[helme1997a] Arne Helme. A System fo Secure User-controlled Electronic Transactions. PhD thesis, University of Twente, 1997.

[hyden1994a] Eoin Andrew Hyden. Operating System Support for Quality of Service. PhD thesis, Wolfson College, University of Cambridge, February 1994.

[johansen1993c] Dag Johansen. The StormCast Applied Approach to Distributed Computing. Dr. sci. thesis, Department of Computer Science, Institute of Mathematical and Physical Sciences, University if Tromsø, May 1993.

[johansen1993d] Dag Johansen. The StormCast Applied Approach to Distributed Computing — Addendum. Dr. sci. thesis, Department of Computer Science, Institute of Mathematical and Physical Sciences, University if Tromsø, May 1993.

[johansen1994a] Stig Johansen. Abstraksjon av kontinuerlige media. Cand. scient. oppgave i informatikk, Seksjon for Informatikk, Institutt for Matematiske Realfag, Universitetet i Tromsø, June 1994.

[karlsen1995a] Randi Karlsen. Flexible Transaction Managment in Federated Database Systems. Dr. scient. thesis, Department of Computer Science, Institute of Mathematical and Physical Sciences, University of Tromsø, January 1995.

[khalaila1997b] Ahmad H. Khalaila. Partial Evaluation and Early Delivery: The Design of a Distributed Pipelined Bushy Query Processing System. Dr. scient. thesis in computer science, Department of Computer Science, Institute of Mathematical and Physical Sciences, University of Tromsø, 1997.

[kvarnrud1991a] Pål Kværnrud. Kjernelaster — en bru mellom arbeidsstasjoner og Mvme147 kort. Diplomoppgave, Universitetet i Tromsø, December 1991.

[lein-mathisen1992a] Terje Lein-Mathisen. A protocol for lock based cache consistensy in the Pegasus file system. Diploma thesis, Department of Computer Science, University of Tromsø / Faculty of Computer Science, University of Twente, The Netherlands, December 1992.

[long1992a] Philip M. Long. Towards a More Comprehensive Theory of Learning in Computers. PhD thesis, University of California, June 1992.

[lopes1997b] Cristina Videira Lopes. D: A Language Framework for Distributed Programming. PhD thesis, College of Computer Science, Northeastern University, November 1997.

[mehus1997a] Sindre Mehus. Type checking and binding of stream interfaces in a multimedia database system. Cand. scient. thesis in computer science, Department of Computer Science, Faculty of Science, University of Tromsø, December 1997.

[nakamura1993a] Akira Nakamura. An Investigation of Real-Time Synchronisation. PhD thesis, Wolfson College, University of Cambridge, Computer Laboratory, December 1993.

[nguyet1991a] Lu Thi Thu Nguyet. Utvikling av et datainnsamlings/presentasjonssystem i sammenheng med StormCast. Semesteroppgave, Universitetet i Tromsø, May 1991.

[nguyet1991b] Lu Thi Thu Nguyet. StormLog — utvikling og realisering av en datalogger for StormCast. Diplomoppgave, Universitetet i Tromsø, December 1991.

[overrein1998a] Øyvind Overrein. Characterization of Vessels and Vessel Generated Surface and Internal Waves Using Multistatic Multifrequency Radar System. Doctoral dissertation, Department of Physics, Faculty of Science, University of Tromsø, 1998.

[papathomas1992a] Michael Papathomas. Language Design Rationale and Semantic Framework for Concurrent Object-Oriented Programming. PhD thesis, Department of Computer Science, University of Geneva, 1992.

[plagemann1996a] Thomas Peter Plagemann. A Framework for Dynamic Protocol Configuration. Dr. thesis, Computer Engineering and Networks Laboratory, Swiss Federal Institute of Technology Zurich, 1996.

[roe1997a] Michael Roe. Cryptography and Evidence. PhD thesis, Clare College, University of Cambridge, 1997.

[satran1991a] Asbjørn Sætran. Monitoring network activity; an application of the COST 11 ter MANDIS approach to monitoring. Cand. scient thesis, University of Tromsø, May 1991.

[smith1982a] Brian Cantwell Smith. Procedural Reflection in Programming Languages. PhD thesis, Massachusetts Institute of Technology, 1982.

[solheim1991a] Inger Solheim. Backpropagation neural networks for facial recognition. Diploma thesis, University of Tromsø, Norway and Los Alamos National Laboratory, Los Alamos, NM, USA, December 1991.

[solheim1998a] Inger Solheim. Bidirectional and Spectral Reflectance Properties of the Cladina stellaris and Flavocetraria nivalis lichens and the Racomitrium lanuginosum moss. Dr. scient. thesis, Department of Physics, Faculty of Science, University of Tromsø, NORUT IT, Tromsø, Norway, May 1998.

[sorensen1993a] Tryggve Sørensen. Design and implementation of MacroScope file system. Master’s thesis, Department of Computer Science, Institute of Mathematical and Physical Sciences, University of Tromsø, 1993.

[stabell-kulo1992a] Tage Stabell-Kulø. En partiell implementasjon av MacroScope*. Cand. scient oppgave, Universitetet i Tromsø, August 1992.

[steen1989a] Roar Steen. Samarbeid mellom regelbaserte systemer. Cand. mag. oppgave, Universitetet i Tromsø, 1989.

[stephenson1991a] Patrick Stephenson. Fast Ordered Multicast. PhD thesis, Cornell University, May 1991.

[sundsfjord1994a] Hans Sundsfjord. Et interaktivt funksjonelt programmeringssystem med eksterne abstrakte datatyper. Hovedfagsoppgave, Seksjon for Informatikk, Institutt for Matematiske Realfag, Universitetet i Tromsø, November 1994.

[vinter1999a] Brian Vinter. PastSet, A Structured Distributed Shared Memory System. Dr. thesis, Department of Computer Science, Faculty of Science, University of Tromsø, June 1999.